OpenStudy (anonymous):

A cell phone plan has a monthly cost that is shown in the table below. What is the correct statement regarding the average rate of change during the 40 minute time of talk? Total minutes of talk time Monthly cost of cell phone 0 $9.95 10 $10.65 20 $11.35 30 $12.05 40 $12.75

OpenStudy (anonymous):

The average rate of change is $1.30, meaning that for each minute of talk time, the monthly bill increases by $1.30. The average rate of change is $0.07, meaning that for each minute of talk time, the monthly bill increases by $0.07. The average rate of change is $1.30, meaning that for every ten minutes of talk time, the monthly bill increases by $1.30. The average rate of change is $0.07, meaning that for every ten minutes of talk time, the monthly bill increases by $0.07.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

0 minutes - $9.95 .... 40 minutes - $12.75 Average rate of change = (12.75 - 9.95) / (40 - 0) = ? dollars per minute of talk time.

OpenStudy (anonymous):

what choice is it ?

OpenStudy (ranga):

What do you get when you plug the numbers into your calculator? (12.75 - 9.95) / (40 - 0) = ?

OpenStudy (anonymous):

0.07

OpenStudy (ranga):

Yes. The answer is B.

OpenStudy (anonymous):

thank you